# A Personalised Formal Verification Framework for Monitoring Activities of Daily Living of Older Adults Living Independently in Their Homes

### v1

@@ 4.3 Formalism and Property Specification
We are interested in modelling behaviour over time. This makes the temporal logic LTL apt for our work as it allows us to specify properties using various temporal operators. In our work, we use LTL extended with past-time operators [27], henceforward referred as past Linear Temporal Logic (pLTL), as it enables reasoning about events that have already occurred. This extension offers practical expressiveness, allowing properties based on sequencing of events to be specified more clearly and compactly.
@@ 4.4 Behaviour-related Properties for Participants
A daily routine corresponds to a subtrace delimited by time markers startActivity and endActivity. To improve clarity and reduce repetition in the specifications, we define the following pattern:
@@
This specifies that participant A moves from the bedroom to the bathroom to take a shower. Since their preference is to do this after waking up, we check this property in accordance to their morning behaviour (see Section 5). The property falls under the general category of the qualitative analysis. For the participant, performing this ADL involves moving from the bedroom to the corridor, then to the bathroom and closing the shower cubicle door. The property (1) specifies that motion in the bedroom () holds until there is motion in the corridor (), which in turn holds until there is a change in the bathroom (). The change in the bathroom must persist until the shower cubicle door is closed (). The bedroom and the corridor each have a single motion sensor; in both cases, one sensor was sufficient to cover the entire area, given the layout. The bathroom has a motion sensor and a contact sensor. A change in the bathroom can be triggered by motion detected in the bathroom () or by closing () or opening () the shower cubicle door.
@@
This specifies that participant A opens the refrigerator to get the salad to feed their pets (before interacting with any other cupboards). Note that since their preference is to do this before breakfast, we check this property in accordance to their morning behaviour (see Section 5). From the qualitative analysis this is an individual-specific property. The property (2) specifies that if the refrigerator door is open () and the door has not been opened in any previous steps (), then none of the drawers or cupboards have been previously opened or closed (), where KitchenCupboards represents all drawers and cupboards in the kitchen.
@@
We also encoded properties specifying the allowable sequence of events imposed by physical constraints. This was done to ensure the sensor events align with the physical layout, thereby verifying correct sensor functioning. We focus here on the physical layout for Participant A’s home (see Figure 2). Property (4) specifies that if motion is detected in the bedroom and, at a later time, in the bathroom, motion must also be detected in the corridor in between.
@@ 5 Experiments, Results and Analysis
From the qualitative analysis, we identify that participant A performs their morning activities around AM every morning. To compensate for early and late mornings, we consider all data captured between AM and AM when evaluating the properties. Similarly, participant B performs their activities around AM every morning; we consider all data captured between AM and AM.
@@
Property 1 was satisfied on all dates between AM and AM except August . Analysis of the counterexample revealed that participant A followed the expected sequence of events between AM and AM, but once in the bathroom, returned to the corridor before going back to the bathroom and closing the shower cubicle door. This unexpected transition caused the property to be evaluated as FALSE, even though it is likely that the activity of taking a shower in the morning was completed. This suggests that temporal sequences may need refinement to accommodate minor deviations in daily routines.
Similarly, Property 2 was violated only on Sunday, August, when participant A did not open the refrigerator before AM. On the same day, a slight delay in the shower routine was observed, indicating that weekend routines could shift in terms of when they occur. This suggests that defining the start of the day relative to the participant’s first set of sensor activations is an aspect worth exploring further. In addition a higher level of specificity may be needed when conducting interviews to address such nuances.
@@
For participant B, we evaluated the property that the medication dispenser was accessed while the participant moved between the kitchen and the lounge between and . Table 4 shows the dates and time intervals for which the property was satisfied. The property was not satisfied on Friday November ; participant B entered the lounge, a necessary step to reach the medication, only after AM. While some ADLs allow for flexibility in when they occur, the intake of medication is a more rigid activity. A deviation in this activity potentially implies a health safety risk.
@@ 6 Conclusion
In this paper, we have presented a framework for representing and reasoning about ADLs performed by older adults. As part of our work, we recruited participants aged to . For each participant, we collected both qualitative and quantitative data over a period ranging from one to three weeks. Our approach leverages contextual information from the participants and data captured by commercially available sensors installed in their homes to build tailored models. We use the qualitative data to specify properties that are tailored to the participants’ preferences and needs. Properties are encoded using LTL extended with past-time operators. Models are automatically generated for each participant and the model checker NuSMV is used to evaluate whether the observed behaviour of participants conform to the encoded properties. To demonstrate the applicability of our framework, we conducted experiments on real-world datasets from two participant living independently. The results support the feasibility of our framework for analysing tailored properties and identifying deviations from expected behaviour. From the case studies presented, we observe that one deviation corresponds to changes in a participant’s routine when taking a shower. Another deviation was triggered because the activity (medication intake) was not performed. Further analysis reveals that, although the activities are not performed at the expected time, they are carried out correctly, albeit a few minutes later. Although this may suggest that the timing for property verification could be relaxed, timely performance of activities such as medication intake might be essential and any violation indicate the need for an intervention.
Acknowledgements
This research was funded by the Legal & General Group (research grant to establish the independent Advanced Care Research Centre at the University of Edinburgh). The funder had no role in conducting the study, interpretation or the decision to submit for publication. The views expressed are those of the authors and not necessarily those of Legal & General.

### v2

@@ 4.3 Formalism and Property Specification
Our focus is on modelling behaviour over time. This makes the temporal logic LTL apt for our work as it allows us to specify properties using various temporal operators. In our work, we use LTL extended with past-time operators [37], henceforward referred to as past Linear Temporal Logic (pLTL), as it enables reasoning about events that have already occurred. This extension offers practical expressiveness, allowing properties based on sequencing of events to be specified more clearly and compactly.
@@ 4.4 Behaviour-related Properties for Participants
A daily routine corresponds to a subtrace delimited by time markers and . To improve clarity and reduce repetition in the specifications, the following pattern is defined:
@@
This specifies that participant A moves from the bedroom to the bathroom to take a shower. Since their preference is to do this after waking up, we check this property in accordance to their morning behaviour (see Section 5). The property falls under the general category of the qualitative analysis. For the participant, performing this ADL involves moving from the bedroom to the hallway, then to the bathroom and closing the shower cubicle door. The property (1) specifies that motion in the bedroom () holds until there is motion in the hallway (), which in turn holds until there is a change in the bathroom (). The change in the bathroom must persist until the shower cubicle door is closed (). The bedroom and the hallway each have a single motion sensor; in both cases, one sensor was sufficient to cover the entire area, given the layout. The bathroom has a motion sensor and a contact sensor. A change in the bathroom can be triggered by motion detected in the bathroom () or by closing () or opening () the shower cubicle door.
@@
This specifies that participant A opens the refrigerator to get the salad to feed their pets (before interacting with any other cupboards). Note that since their preference is to do this before breakfast, we check this property in accordance to their morning behaviour (see Section 5). From the qualitative analysis this is an individual-specific property. The property (2) specifies that if the refrigerator door is open () and the door has not been opened in any previous steps (), then none of the drawers or cupboards have been previously opened or closed (), where represents all drawers and cupboards in the kitchen.
@@
Properties specifying the allowable sequence of events imposed by physical constraints were also encoded. This was done to ensure the sensor events align with the physical layout, thereby verifying correct sensor functioning. We focus here on the physical layout for Participant A’s home (see Figure 2). Property (4) specifies that if motion is detected in the bedroom and, at a later time, in the bathroom, motion must also be detected in the hallway in between.
@@ 5 Experiments, Results and Analysis
From the qualitative analysis, it is identified that participant A performs their morning activities around AM every morning. To compensate for early and late mornings, we consider all data captured between AM and AM when evaluating the properties. Similarly, participant B performs their activities around AM every morning; we consider all data captured between AM and AM.
@@
Property 1 was satisfied on all dates between AM and AM except August . Analysis of the counterexample revealed that participant A followed the expected sequence of events between AM and AM, but once in the bathroom, returned to the hallway before going back to the bathroom and closing the shower cubicle door. This unexpected transition caused the property to be evaluated as FALSE, even though a morning shower activity was almost certainly completed. This indicates that the apparent violation arose from a minor deviation in the sequence, suggesting that temporal requirements may need refinement to accommodate minor deviations in daily routines.
Similarly, Property 2 was violated only on Sunday, August, when participant A did not open the refrigerator before AM. On the same day, a slight delay in the shower routine was observed, indicating that weekend routines could be shifted to later times. This suggests that defining the start of the day relative to the participant’s first set of sensor activations is an aspect to be explored further. In addition a higher level of specificity may be needed when conducting interviews to address such nuances.
@@
For participant B, we evaluated whether the scheduled medication intake (see Property 3) occurred while the participant moved between the kitchen and the lounge between and . Table 4 shows the dates and time intervals for which the property was satisfied. The property was not satisfied on Friday November ; participant B entered the lounge, a necessary step to reach the medication, only after AM. While some ADLs allow for flexibility in when they occur, scheduled medication intake is a more rigid requirement. A deviation in this activity could indicate a health safety risk.
@@
Limitations
This study has some limitations. Sensor placement was sometimes restricted by participant preferences, while pets and unexpected visitors occasionally introduced noise, as the models assume a single participant. The sensors provide only evidence of events (e.g. refrigerator opening) and do not guarantee inherent interpretation (e.g. food intake). Data collection was constrained by study duration, resources and participant availability; consequently, the study included a small number of participants. Malfunctions were observed in three motion sensors (about 1% of the data), which were excluded from analysis. Importantly, though, our person-centred approach mitigates challenges, such as sensor placement, ensuring relevance to participants’ salient ADLs.
@@ 6 Conclusion
In this paper, we have presented a framework for representing and reasoning about ADLs performed by OAs. As part of our work, we recruited participants aged to . For each participant, we collected both qualitative and quantitative data over a period ranging from one to three weeks. Our approach leverages contextual information from the participants and data captured by commercially available sensors installed in their homes to build tailored models. We specify properties tailored to the needs and preferences expressed by the participants and encode them in LTL extended with past-time operators. Models are automatically generated for each participant and the model checker NuSMV is used to evaluate whether the observed behaviour of participants conforms to the encoded properties.
To demonstrate the applicability of our framework, we conducted experiments on real-world datasets from two participant living independently. The results support the feasibility of our framework for analysing tailored properties and identifying deviations from expected behaviour. From the case studies presented, we observe that one deviation corresponds to changes in a participant’s routine when taking a shower. Another deviation was triggered because the activity (medication intake) was not performed. Further analysis reveals that, although the activities are not performed at the expected time, they are carried out correctly outside the anticipated time period. Although this may suggest that the timing for property verification could be relaxed, timely performance of activities such as scheduled medication intake might be essential and a violation may indicate the need for an intervention. Overall, our initial findings show the potential of the framework to be used in the home of OAs living independently. However, even in specific settings where an individual’s context and their preferences and needs are taken into account, ADL monitoring remains challenging. This is because human behaviour and sensing granularity can result in ambiguous interpretations, such as the act of opening a refrigerator not necessarily indicating food consumption.